What is the value of q in this expression when p equals 4?
p squared - qp - 12=0
A. 1
B. -1
C. 2
D. -2
E. 3

Answer by themathtutor2009(81)

p^2-qp-12=0

4^2-q(4)-12=0

16-4q-12=0

4-4q=0

-4q=-4

q=(-4)/(-4)

q=1